# Quantum Generative Diffusion Model: A Fully Quantum-Mechanical Model for Generating Quantum State Ensemble

### v2

@@ Abstract
Classical diffusion models have shown superior generative results and have been applied to many problems. Exploring these models in the quantum domain can advance the field of quantum generative learning. In this paper, we introduce the Quantum Generative Diffusion Model (QGDM), a simple and elegant quantum counterpart of classical diffusion models.
The core idea of QGDM is that any target quantum state can be transformed into a completely mixed state, which has the highest entropy and maximum uncertainty about the system, through a non-unitary forward process. Subsequently, a trainable backward process can be used to recover the target state from the completely mixed state. The design requirements for QGDM’s backward process include ensuring non-unitarity while maintaining a low number of parameters. To achieve this, we introduce partial trace operations in the backward process to enforce non-unitary. Additionally, we control the number of trainable parameters by using a parameter-sharing strategy and incorporating temporal information as an input in the backward process. Furthermore, we introduce a resource-efficient version of QGDM, which reduces the number of auxiliary qubits while preserving impressive generative capabilities.
Our proposed models exhibit better convergence performance than Quantum Generative Adversarial Networks (QGANs) because our models optimize a convex distance function using gradient descent. Comparative results with QGANs demonstrate the effectiveness of our models in generating both pure and mixed quantum states. Notably, our models achieve 53.03% higher fidelity in mixed-state generation tasks compared to QGANs. These results highlight the potential of the proposed models to tackle challenging quantum generation tasks.
@@
I Introduction
Figure 1: Visualizing the core idea of the proposed models in the Bloch sphere.
@@ I Introduction
Generative models have attracted widespread attention and achieved significant success in academic and industrial fields [1, 2]. Notably, in the past two years, large models such as ChatGPT and Stable Diffusion have shown remarkable performance and creativity in the fields of human-machine dialogue and image generation [3, 4]. These models have made technological breakthroughs and significantly impacted daily life [5]. The core goal of generative models is to learn the underlying probability distributions from a given dataset. Once trained, these models can generate new data that follow the distribution, demonstrating an understanding of the underlying patterns in the data.
@@
Quantum generative models [6], as the quantum counterparts of classical generative models in the field of quantum machine learning [7, 8, 9, 10, 11], have sparked great interest in the academic community. Quantum generative models merge the advantages of quantum computing and classical generative models, offering new possibilities for data generation. There have been several preliminary explorations of various quantum generative models, including the Quantum Generative Adversarial Network (QGAN) [12, 13, 14], the Quantum Circuit Born Machine (QCBM) [15, 16], the Quantum Variational Autoencoder (QVAE) [17], and the Quantum Boltzmann Machine (QBM) [18]. Among them, QGAN, the quantum counterpart of the classical Generative Adversarial Network (GAN) [19], has received considerable attention for its remarkable applications, including, but not limited to, generating quantum state [13, 20, 21], image generation [22, 23, 24], and fitting classical discrete distributions [14, 25, 26]. QGAN provides a pathway for researchers to more effectively simulate and explore complex quantum systems, especially in generating and processing data related to quantum states.
I-A Motivations
@@
Although QGANs are applied in various fields, they inherit challenges in convergence from classical Generative Adversarial Networks. QGANs may experience model collapse, leading to oscillations in the loss of the generator and discriminator without converging to a global optimum [21]. Furthermore, QGANs face a unique challenge in the quantum domain as they cannot achieve convergence in generating mixed states [27]. Specifically, according to Helstrom’s measurement theory, the analytic optimal output of the generator under a minimax optimization strategy is a pure state, which demonstrates its inability to produce a target mixed state. The aforementioned challenges make the training of QGANs difficult and limit their broader application in critical quantum state generation tasks. These limitations highlight the need for developing more effective alternatives to address these complex tasks of quantum state generation.
@@
To achieve this goal, it is important to explore the quantum analogs of classical generative diffusion models [28, 29, 30, 31]. Classical diffusion models have been proven to have superior generative performance in many fields compared to GAN, and they have more friendly convergence properties than classical GANs [32, 33]. Quantum analogs of diffusion models may offer similar benefits by setting the objective function to minimize a convex distance function between the target quantum state and the generated quantum state. Optimizing a convex function using gradient descent typically results in better convergence properties [34]. Therefore, the development of such models is expected to bypass some challenges faced by QGANs and contribute to the field of quantum generative learning. These considerations motivated us to undertake this work.
I-B Contributions
@@
Inspired by classical diffusion models, this paper introduces a novel quantum generative model known as the Quantum Generative Diffusion Model (QGDM) for quantum state generation tasks. The core idea of QGDM is shown in Fig. 1. Any target quantum state ($\rho_{0}$) can be transformed into a completely mixed state ($\rho_{T}$) through a forward (diffusion) process. Subsequently, a trainable backward (denoising) process can be used to recover the target state from the completely mixed state. Our objectives are to define a timestep-dependent forward process, design an effective quantum state backward process, and benchmark our model in comparison to other quantum generative models in pure and mixed state generation tasks. Our contributions are summarized as follows:
@@
    We propose an elegant quantum version of the diffusion model called the Quantum Generative Diffusion Model (QGDM). The core idea of QGDM is that any target quantum state can be diffused into a completely mixed state by using a forward process. Then, using a learnable backward process, the original target quantum state can be recovered from the completely mixed state. The forward process employs a non-unitary depolarization channel model. To reverse the diffusion process, we introduce a partial trace operation during denoising, which traces out information from the subsystem and makes the backward process non-unitary. Additionally, we use a parameter-sharing strategy and include temporal information as an inputs, allowing the backward processes at all timesteps to share the same parameters.
@@
    We introduce a resource-efficient version of QGDM (RE-QGDM) to reduce the number of auxiliary qubits in the backward process, thereby enhancing the model’s efficacy in addressing more complex and large-scale problems.
@@
    Numerical simulation results demonstrate that the proposed models outperform QGANs in generating both mixed states and pure states. Notably, in generating mixed states, QGDM achieves 22.17% higher fidelity compared to QGANs for tasks involving 1 to 4 qubits, while RE-QGDM achieves 64.66% higher fidelity compared to QGANs for tasks involving 1 to 8 qubits.
The remainder of this article is structured as follows. Initially, in Section II, we delve into the existing literature and the foundational work relevant to quantum generative models. Subsequently, in Section III, we present an in-depth exposition of our innovative Quantum Generative Diffusion Model (QGDM) framework. Further, Section IV is dedicated to introducing a Resource-Efficient iteration of the QGDM (RE-QGDM). Section V analyzes the rationale behind the design of the backward process and presents simulation results to support these insights. These analyses and simulation outcomes are poised to aid future research in similar models. The numerical simulations setup and comparative results with other quantum generative models are shown in Section VI. Finally, Section VIII concludes this article. The notations are summarized in Table I.
@@
$t$ Timestep
$T$ Total Timesteps of the Forward Process
$|\psi\rangle$ Quantum State
$\rho$ Density Matrix of A Quantum System
@@
$\rho_{T}$ Completely Mixed State
$\tau_{t}$ Timestep Embedding Quantum State
$\mathbb{I}$ Identity Matrix
$\alpha_{t}$ Noise Parameter Dependent on $t$
@@
$N_{\text{time}}$ Number of Qubits in the Timestep Embedding Circuit
@@
$\mathcal{L}$ Total Loss Function
II Related Works
II-A Preliminary
We first introduce the necessary quantum information backgrounds. For more details, please refer to Ref. [35].
II-A1 Pure Quantum State
In quantum computing, a state of an $N$-qubit can be represented by a vector in a complex Hilbert space $\mathbb{C}^{d}$ with unit length, denoted as $|\psi\rangle\in\mathbb{C}^{d}$, where the ket notation |$\rangle$| ⟩ denotes a column vector and $d=2^{N}$ is the dimension of Hilbert space. The bra notation $\langle\psi|=|\psi\rangle^{\dagger}$ represents a row vector, with $\dagger$ denoting the conjugate transpose. The evolution of a pure state can be represented by applying a quantum circuit or a quantum gate: $|\psi^{\prime}\rangle=U|\psi\rangle$, where $U$ is a unitary operator satisfying $UU^{\dagger}=\mathbb{I}$, and $\mathbb{I}$ is the identity matrix.
II-A2 Mixed Quantum State
A mixed state of a quantum system is a probabilistic mixture of pure states. It can be represented by a density matrix (or density operator): $\rho=\sum_{i}p_{i}|\psi_{i}\rangle\langle\psi_{i}|$, where the ensemble of pure states $\{p_{i},|\psi_{i}\rangle\}$ indicating that the system is in state $|\psi_{i}\rangle$ with a probability of $p_{i}$. The density matrix is particularly useful in contexts where the exact state of the system is not known or when considering subsystems of entangled pairs. The evolution of a mixed state is described by $\rho^{\prime}=U\rho U^{\dagger}$.
II-A3 Tensor Product
@@ II-A4 Partial Trace
Suppose we have two quantum systems, A and B. Their composite system state is represented by the density matrix $\rho_{AB}$. Partial trace operation yields a reduced density matrix for subsystem A, calculated by: $\rho_{A}=\text{tr}_{B}(\rho_{AB})$.
II-B Quantum Generative Models
@@
Quantum generative models, a pivotal sector in quantum machine learning, encompass various models including Quantum Circuit Born Machine (QCBM) [15], Quantum Boltzmann Machine (QBM) [18], Quantum Variational Autoencoder (QVAE) [17], and Quantum Generative Adversarial Network (QGAN) [12, 13].
@@
The Quantum Circuit Born Machine (QCBM) uses quantum circuits to generate a classical bit string from specific probability distributions, a concept first proposed by Benedetti et al. [15]. This model has seen substantial advances [16, 36, 37, 38], finding applications in various fields, including Monte Carlo simulations[39], financial data generation[40, 41], and joint distribution learning[40, 42]. The Quantum Boltzmann Machine (QBM), functioning as a quantum counterpart of the classical Boltzmann Machine [43, 44], uses a quantum device to prepare Boltzmann distributions, by replacing the units in the classical Boltzmann Machine with qubits and substituting the energy function with a parameterized Hamiltonian. QBM was first demonstrated by Amin et al. [18], leading to significant model enhancements [45, 46, 47]. Khoshaman et al. [17] introduced the Quantum Variational Autoencoder (QVAE), using the QBM as a prior in Discrete VAE [48], with subsequent developments including annealer-based [49] and gate-based QVAE [50]. In the realm of QGANs, Lloyd et al. [12] developed a comprehensive taxonomy, classifying them by the quantum or classical nature of their components. This was empirically supported by Dallaire-Demers and Killoran [13], who demonstrated the efficacy of QGAN in generating the desired quantum data. Since then, the field has continuously explored applications in quantum state generation [51, 20, 27, 21], image generation [22, 23, 24], and discrete data generation[14, 25, 26]. For an in-depth review of these quantum generative models, please refer to the comprehensive work by Tian et al. [6].
@@
During our research, several papers [52, 53, 54] also proposed quantum analogs to the classical diffusion model. Ref. [52] introduces several strategies and theoretical tools that could potentially be applied to develop quantum versions of diffusion models. However, the paper lacks specific implementation examples and experimental validation. In contrast, our work introduces a detailed concept and delves into technical details, including comparisons with other quantum generative models. Ref.[53] proposes the QuDDPM algorithm to generate individual pure states from a target distribution. Instead, our work aims to generate the average state of the distribution, which is also an objective considered in other research on quantum generative models [13, 20, 27, 21]. Additionally, the backward process in the QuDDPM algorithm uses a different set of parameters at each timestep. In contrast, we use a parameter-sharing strategy, ensuring that the same set of parameters is used for the backward process at every timestep. This approach keeps the total number of parameters constant, regardless of the number of timesteps, thereby avoiding an increase in parameters as the number of timesteps grows. Ref. [54] uses a classical approach for their diffusion process, while the denoising process is performed using a trainable quantum circuit for the denoising task. However, our method is fully quantum mechanical, encompassing both diffusion and denoising processes.
III Quantum Generative Diffusion Model
III-A General Framework
@@
Fig. 2 shows a general framework of the proposed Quantum Generative Diffusion Model (QGDM). QGDM includes a forward process (or diffusion process) $\mathcal{E}(\rho_{t-1},t)$ and a backward process (or denoising process) $f_{\boldsymbol{\Theta}}(\rho_{t},t)$. The diffusion process $\mathcal{E}(\rho_{t-1},t)$ takes an $N$-qubit quantum state $\rho_{t-1}$ and a timestep scalar $t$ as input, producing a quantum state $\rho_{t}$ with added noise. Similar to classical diffusion models, in QGDM, a “timestep” refers to a discrete step in the model’s processes where noise is sequentially added to (forward process) or subtracted from a quantum state (backward process). In this paper, $\mathcal{E}(\rho_{t-1},t)$ is modeled as a depolarizing channel. When $t=0$, $\rho_{0}$ represents the target quantum state. In contrast, at $t=T$, $\rho_{T}$ is a completely mixed state. The denoising process employs a model $f_{\boldsymbol{\Theta}}(\rho_{t},t)$ with trainable parameters $\boldsymbol{\Theta}$, which denoises the noisy input quantum state $\rho_{t}$, producing a cleaner quantum state $\rho_{t-1}$. After applying the function $f_{\boldsymbol{\Theta}}(\cdot,\cdot)$ $T$ times to $\rho_{T}$, the target quantum state $\rho_{0}$ is recovered.
@@
Remark 1: Here we discuss the difference between the notations used in this paper and those used in classical diffusion model papers [29, 32, 33]. In quantum information, the quantum operation formalism $\rho^{\prime}=\mathcal{E}(\rho)$ is a powerful tool for describing the evolution of quantum states [35], where $\mathcal{E}$ represents a quantum operation, and $\rho$ and $\rho^{\prime}$ are the initial and final states, respectively. This formalism is especially useful for describing open quantum systems or the stochastic evolution of quantum states. In this paper, we utilize the quantum operation formalism, denoting forward and backward processes as $\mathcal{E}(\rho_{t-1},t)$ and $f_{\boldsymbol{\Theta}}(\rho_{t},t)$, respectively. These notations do not explicitly include the output term. In contrast, in classical diffusion model papers, forward and reverse distributions are defined as $q(x_{t}|x_{t-1})$ and $p_{\boldsymbol{\theta}}(x_{t-1}|x_{t})$. These notations explicitly include the output term ($i.e.$, $x_{t}$ in $q(x_{t}|x_{t-1})$ and $x_{t-1}$ in $p_{\boldsymbol{\theta}}(x_{t-1}|x_{t})$). The difference between the notations in our paper and those in classical diffusion model papers can be summarized as follows. In the classical case, the evolution process of a sample ($x_{t-1}$ or $x_{t}$) is probabilistic. Therefore, the forward or backward process uses conditional probability notation. In the quantum case, the system’s state is described by a density matrix. The evolution of the density matrix is deterministic, but the density matrix itself represents a probabilistic quantum system. Thus, we use the quantum operation formalism to describe the forward and backward processes. This distinction between quantum and classical systems is emphasized to aid readers who may not be familiar with quantum computation in understanding our notations.
III-B Forward Process

### v3

@@ Abstract
Classical diffusion models have shown superior generative results. Exploring them in the quantum domain can advance the field of quantum generative learning. This work introduces Quantum Generative Diffusion Model (QGDM) as their simple and elegant quantum counterpart. Through a non-unitary forward process, any target quantum state can be transformed into a completely mixed state that has the highest entropy and maximum uncertainty about the system. A trainable backward process is used to recover the former from the latter. The design requirements for its backward process includes non-unitarity and small parameter count. We introduce partial trace operations to enforce non-unitary and reduce the number of trainable parameters by using a parameter-sharing strategy and incorporating temporal information as an input in the backward process. We present QGDM’s resource-efficient version to reduce auxiliary qubits while preserving generative capabilities. QGDM exhibits faster convergence than Quantum Generative Adversarial Network (QGAN) because its adopted convex-based optimization can result in faster convergence. The results of comparing it with QGAN demonstrate its effectiveness in generating both pure and mixed quantum states. It can achieve 53.02% higher fidelity in mixed-state generation than QGAN. The results highlight its great potential to tackle challenging quantum generation tasks.
@@
1 Introduction
@@
Generative models have attracted widespread attention and achieved significant success in industrial fields [1]. Notably, in the past two years, large models such as ChatGPT and Stable Diffusion showed remarkable performance and creativity in human-machine dialogue and image generation [2, 3]. They have made technological breakthroughs and significantly impacted our daily life. Their essence is to learn the underlying probability distributions from a given dataset. Once trained, these models can generate new data that follow the distribution, demonstrating an understanding of the underlying patterns in the data.
@@
Quantum generative models [4], as the quantum counterparts of classical generative models in quantum machine learning [5, 6, 7], have sparked great interest in the research community. They merge the advantages of quantum computing and classical generative models, offering new possibilities for data generation. There have been several explorations of such models, including Quantum Generative Adversarial Network (QGAN) [8, 9], Quantum Circuit Born Machine (QCBM) [10, 11], Quantum Variational Autoencoder (QVAE) [12], and Quantum Boltzmann Machine (QBM) [13]. Among them, as the quantum counterpart of the classical Generative Adversarial Network (GAN) [14, 15, 16, 17, 18], QGAN has received considerable attention for its remarkable applications, $e.g.$, generating quantum states [9, 19, 20, 21], image generation [22, 23, 24, 25], and fitting classical distributions [26, 27, 28, 29, 30]. It provides a pathway for researchers to more effectively simulate and explore complex quantum systems, especially in generating and processing data related to quantum states.
1.1 Motivations
@@
Despite their various applications, QGANs inherit challenges in convergence from GANs. QGANs may experience model collapse, leading to oscillations in the loss of the generator and discriminator without converging to a global optimum [21]. Furthermore, they face a unique challenge in the quantum domain as they cannot achieve convergence in generating mixed states [20]. Specifically, according to Helstrom’s measurement theory, the analytic optimal output of the generator under a minimax optimization strategy is a pure state, which demonstrates its inability to produce a target mixed state. The aforementioned challenges make their training difficult and limit their broader application in critical quantum state generation tasks. These limitations highlight the need for developing more effective alternatives to perform the complex tasks of quantum state generation.
Figure 1: The core idea of the proposed QGDM.
@@
To do so, it is important to explore the quantum analogs of classical generative diffusion models [31, 32, 33, 34]. These models have been proven to have superior generative performance in many fields over GANs with better convergence [35, 36]. Their quantum analogs may offer similar benefits. For example, if we can set up a convex objective function that minimizes the distance between the target and generated quantum states, then using gradient descent to optimize the convex function can result in better convergence properties [37]. Therefore, the development of such models is expected to address the challenges faced by QGANs and make a breakthrough contribution to the field of quantum generative learning. These considerations have motivated us to undertake this work.
1.2 Contributions
@@
Inspired by classical diffusion models, this paper introduces a novel quantum generative model known as QGDM for quantum state generation tasks. Fig. 1 shows its core idea in a Bloch sphere: any target quantum state ($\rho_{0}$) can be transformed into a completely mixed state ($\rho_{T}$) through a forward (diffusion) process; subsequently, a trainable backward (denoising) process can be used to recover the former from the latter. Our objectives are to a) define a timestep-dependent forward process, b) design an effective quantum state backward process, and c) benchmark our model against other models in pure and mixed state generation tasks. We aim to make the following novel and significant contributions to the field of quantum generative learning:
@@
    We propose an elegant quantum version of diffusion model called QGDM. Any target quantum state can be diffused into a completely mixed state by using a forward process. Then, using a learnable backward process, the former can be recovered from the latter. The forward process employs a non-unitary depolarization channel model. To reverse the diffusion process, we introduce a partial trace operation during denoising, which traces out information from the subsystem and makes the backward process non-unitary. Additionally, to reduce the number of parameters, we use a parameter-sharing strategy and include temporal information as an input, allowing the backward processes at all timesteps to share the same parameters.
@@
    We introduce a Resource-efficient version of QGDM (RQGDM) to minimize the number of auxiliary qubits required in the backward process. Its core idea is to utilize a parameterized quantum circuit to extract a low-dimensional representation of quantum data from its original high-dimensional Hilbert space. This compression is crucial for the backward process since the low-dimensional representation uses much fewer qubits than the input state. Consequently, RQGDM enhances the model’s ability to tackle large-scale problems efficiently.
@@
    We conduct numerical simulation whose results demonstrate that the proposed models well outperform QGANs in generating both mixed states and pure states. Notably, in generating mixed states, QGDM achieves 22.17% higher fidelity than QGANs for tasks involving 1 to 4 qubits, while RQGDM achieves 64.66% higher one than QGANs for tasks involving 1 to 8 qubits.
The remainder of this article is structured as follows. Section 2 briefs the existing literature and the foundational work relevant to quantum generative models. Section 3 presents an in-depth exposition of QGDM. Section 4 introduces RQGDM. Section 5 analyzes the rationale behind the design of the backward process and presents simulation results. Section 6 gives the numerical simulation setup and comparative results. Finally, Section 7 concludes this article.
2 Preliminary and Related Work
2.1 Preliminary
We first introduce the necessary quantum information background. For more details, please refer to Ref. [38]. The notations are summarized in Table I.
@@ 2.1 Preliminary
$\tau_{t}$ Timestep Embedding Quantum State at Timestep $t$
$\alpha_{t}$ Noise Parameter Dependent at Timestep $t$
@@
$N_{\tau}$ Number of Qubits in the Timestep Embedding Circuit
@@
2.1.1 Pure Quantum State
In quantum computing, a state of an $N$-qubit can be represented by a vector in a complex Hilbert space $\mathbb{C}^{d}$ with unit length, denoted as $|\psi\rangle\in\mathbb{C}^{d}$, where the ket notation |$\rangle$| ⟩ denotes a column vector and $d=2^{N}$ is the dimension of Hilbert space. The bra notation $\langle\psi|=|\psi\rangle^{\dagger}$ represents a row vector, with $\dagger$ denoting the conjugate transpose. The evolution of a pure state can be represented by: $|\psi^{\prime}\rangle=U|\psi\rangle$, where $U$ is a unitary operator satisfying $UU^{\dagger}=\mathbb{I}$, and $\mathbb{I}$ is the identity matrix.
2.1.2 Mixed Quantum State
A mixed state of a quantum system is a probabilistic mixture of pure states. It can be represented by a density matrix (or density operator): $\rho=\sum_{i}p_{i}|\psi_{i}\rangle\langle\psi_{i}|$, where the ensemble of pure states $\{p_{i},|\psi_{i}\rangle\}$ indicates that the system is in state $|\psi_{i}\rangle$ with a probability of $p_{i}$. This matrix is particularly useful in contexts where the exact state of a system is not known or when considering the subsystems of entangled pairs. The evolution of a mixed state is described by $\rho^{\prime}=U\rho U^{\dagger}$.
2.1.3 Tensor Product
@@ 2.1.4 Partial Trace
Suppose that we have two quantum systems, A and B. Their composite system state is represented by the density matrix $\rho_{AB}$. The partial trace operation yields a reduced density matrix for subsystem A, calculated by $\rho_{A}=\text{tr}_{B}(\rho_{AB})$. Its definition is $\text{tr}_{B}\left(|a_{1}\rangle\langle a_{2}|\otimes|b_{1}\rangle\langle b_{2}|\right)=|a_{1}\rangle\langle a_{2}|\text{tr}\left(|b_{1}\rangle\langle b_{2}|\right)$, where $|a_{1}\rangle$ and $|a_{2}\rangle$ are any two vectors in A, while $|b_{1}\rangle$ and $|b_{2}\rangle$ are any two vectors in B.
2.2 Quantum Generative Models
@@
Quantum generative models, a pivotal sector in quantum machine learning, encompass various models including QCBM [10], QBM [13], QVAE [12], and QGAN [8, 9].
@@
QCBM uses quantum circuits to generate a classical bit string from specific probability distributions [10]. This model has substantial advanced [11, 39, 40, 41], finding its applications in various fields, $e.g.$, Monte Carlo simulations[42], financial data generation[43], and joint distribution learning[43, 44]. QBM functioning as a quantum counterpart of the classical Boltzmann Machine [45, 46], uses a quantum device to prepare Boltzmann distributions, by replacing the units in the Boltzmann Machine with qubits and substituting the energy function with a parameterized Hamiltonian. QBM was first demonstrated in [13], has its enhancements [47, 48]. Khoshaman et al. [12] introduced QVAE, using QBM as a prior in Discrete VAE [49], with developments including annealer-based [50] and gate-based QVAE [51]. In the realm of QGANs, Lloyd et al. [8] developed a comprehensive taxonomy, classifying them by the quantum or classical nature of their components. This was empirically supported by Dallaire-Demers and Killoran [9], who demonstrated the efficacy of QGAN in generating the desired quantum data. Since then, the field has continuously explored applications in quantum state generation [52, 20, 21], image generation [22, 23, 24], and discrete data generation[28, 27, 30]. For an in-depth review of these quantum generative models, please refer to [4].
@@
During our research, several studies [53, 54, 55] have proposed quantum analogs to the classical diffusion model. Parigi et al. [53] introduce several strategies and tools that could be applied to develop quantum versions of diffusion models. However, they fail to offer specific implementation examples and experimental validation. In contrast, our work introduces a detailed concept and delves into technical details, including comparisons with other quantum generative models. Zhang et al. [54] propose quantum denoising diffusion probabilistic model (QuDDPM) to generate individual pure states from a target distribution. Instead, our work aims to generate the average state of the distribution, which is also an objective considered in other research on quantum generative models [9, 20, 21]. Additionally, the backward process in QuDDPM uses a different set of parameters at each timestep. In contrast, we use a parameter-sharing strategy, ensuring that the same set of parameters is used for the backward process at every timestep. This approach keeps the total number of parameters constant, regardless of the number of timesteps, thereby avoiding an increase in parameters as the number of timesteps grows. Cacioppo et al. [55] use a classical approach for their diffusion process, while the denoising process is performed via a trainable quantum circuit for the denoising task. However, our method is fully quantum-mechanical, encompassing both diffusion and denoising processes.
3 Quantum Generative Diffusion Model (QGDM)
3.1 General Framework
@@
Fig. 2 shows a framework of QGDM. It includes a forward process (or diffusion one) $\mathcal{E}(\rho_{t-1},t)$ and a backward process (or denoising one) $f_{\boldsymbol{\Theta}}(\rho_{t},t)$. $\mathcal{E}(\rho_{t-1},t)$ takes an $N$-qubit quantum state $\rho_{t-1}$ and a timestep scalar $t$ as input, producing a quantum state $\rho_{t}$ with added noise. Similar to classical diffusion models, in QGDM, a “timestep” refers to a discrete step in the model’s processes where noise is sequentially added to (forward process) or subtracted from a quantum state (backward process). In this paper, $\mathcal{E}(\rho_{t-1},t)$ is modeled as a depolarizing channel. When $t=0$, $\rho_{0}$ represents the target quantum state. In contrast, at $t=T$, $\rho_{T}$ is a completely mixed state. The denoising process employs a model $f_{\boldsymbol{\Theta}}(\rho_{t},t)$ with trainable parameters $\boldsymbol{\Theta}$, which denoises the noisy input quantum state $\rho_{t}$, producing a cleaner quantum state $\rho_{t-1}$. After applying the function $f_{\boldsymbol{\Theta}}(\cdot,\cdot)$ for $T$ times to $\rho_{T}$, the target quantum state $\rho_{0}$ is recovered.
@@
Remark 1: Let us discuss the difference between the notations used in this paper and those used in classical diffusion model papers [32, 35, 36]. In quantum information, the quantum operation formalism $\rho^{\prime}=\mathcal{E}(\rho)$ is a powerful tool for describing the evolution of quantum states [38], where $\mathcal{E}$ represents a quantum operation, and $\rho$ and $\rho^{\prime}$ are the initial and final states, respectively. This formalism is especially useful for describing open quantum systems or the stochastic evolution of quantum states. In this paper, we utilize the quantum operation formalism, denoting forward and backward processes as $\mathcal{E}(\rho_{t-1},t)$ and $f_{\boldsymbol{\Theta}}(\rho_{t},t)$, respectively. These notations do not explicitly include the output term. In contrast, in classical diffusion models [31, 32], forward and reverse distributions are defined as $q(x_{t}|x_{t-1})$ and $p_{\boldsymbol{\theta}}(x_{t-1}|x_{t})$. These notations explicitly include the output term ($i.e.$, $x_{t}$ in $q(x_{t}|x_{t-1})$ and $x_{t-1}$ in $p_{\boldsymbol{\theta}}(x_{t-1}|x_{t})$). The difference between the notations in our paper and those in classical diffusion model papers can be summarized as follows. In the classical case, the evolution process of a sample ($x_{t-1}$ or $x_{t}$) is probabilistic. Therefore, the forward or backward process uses conditional probability notation. In the quantum case, the system’s state is described by a density matrix. The evolution of the density matrix is deterministic, but the density matrix itself represents a probabilistic quantum system. Thus, we use the quantum operation formalism to describe the forward and backward processes. This distinction between quantum and classical systems aids readers who may not be familiar with quantum computation in understanding our notations.
3.2 Forward Process